OpenStudy (anonymous):

A survey found that 3 out of 100 people have red hair. On the basis of this survey, how many people in a group of 12,000 people are likely to have red hair? A. 36 B. 360 C. 4000 D. 3

OpenStudy (anonymous):

3/100 = x/12000

OpenStudy (pinkcandyrosez882):

Sorry I think my answer was wrong. I agree with kelliegirl33

OpenStudy (anonymous):

now cross multiply and solve for x

OpenStudy (owlcoffee):

Let's just analyze it using the first data. What it basically means is that in a group of a hundred people it's likely to find 3 of them with red hair. Therefore, that means that it is a 3% or if you like: \[\frac{ 3 }{ 100 } = a\] We call "a" the amount of red haired people in any given amount of people, so now we will take a look at the quantity of 12,000, we can calculate it with a rule of three. we enounce it like: "if in a group of a hundred people, 3 are red haired, how many red haired will it be in 12,000?" So we will multiply 12,000 by 3 and divide it by 100 to obtain the value we are looking for: \[\frac{ (12,000)(3) }{ 100 }=\frac{ 36,000 }{ 100 }=360\]

OpenStudy (anonymous):

C

OpenStudy (anonymous):

not C

OpenStudy (anonymous):

owlcoffee just gave you the answer

OpenStudy (anonymous):

plz refrain from giving dirext answers\